A car moves around a bend of radius 20m with an acceleration of 12ms-2. What is the maximum speed the car can attain without skidding?


1
Expert's answer
2021-11-11T12:00:22-0500

The cenripetal acceleration

"a=\frac{v^2}{r}"

Hence, the speed of the car

"v=\sqrt{ar}=\sqrt{12*20}=15.5\:\rm m/s"